Nico is saving money for his college education. He invests some money at 5%, and $1900 less than that amount at 4%. The investments produced a total of $185 interest in 1 yr. How much did he invest at each rate?
Found 2 solutions by checkley77, nyc_function:
Answer by checkley77(12844)

.05x+.04(x-1,900)=185
.05x+.04x-76=185
.09x=185+76
.09x=261
x=261/.09
x=2,900 amount invested @ 5%.
2,900-1,900=1,000 amount invested @ 4%.
Proof:
.05*2,900+.04*1,000=185
145+40=185
185=185

Answer by nyc_function(2741)

Let x be the amount he invests at 5%. Then the amount at 4% will be x - 1900. His interest for one year is
185 = 0.05x + 0.04(x - 1900)
= 0.05x + 0.04x - 76
185 + 76 = 0.09x
x = 261/0.09 = 2900
So, he invested $2900 at 5% and $1000 at 4%.